EXPERIMENTAL INVESTIGATION OF THE DISSOCIATION MECHANISMS OF He⋯ICl(B,v′) VAN DER WAALS COMPLEXES

\underline{JOSHUA P. DARR} and RICHARD A. LOOMIS, Department of Chemistry, Washington University, One Brookings Drive, CB 1134, Saint Louis, MO 63130.

Two-laser, pump-probe spectroscopy experiments reveal different dissociation pathways for He⋯ICl(B,v′) intermolecular levels. Weakly bound He⋯I^{35}Cl(B,v′=3) complexes that are prepared with different quanta of intermolecular vibrational excitation are observed to undergo vibrational predissociation to yield predominantly I^{35}Cl(B,v′=2) products. The I^{35}Cl(B,v′=2) rotational product state distributions formed following the excitation of different intermolecular vibrational levels and rotational states are all bimodal. Less rotational excitation is found in the products when preparing He⋯ICl(B,v′) bending levels that preferentially sample the linear He−I−Cl and anti-linear I−Cl−He orientations than when preparing the lowest intermolecular vibrational energy level, which has a preferred T-shaped geometry. Additionally, linear He⋯I^{35}Cl(B,v′=2) complexes undergo a direct dissociation mechanism when prepared above the He + I^{35}Cl(B,v′=2) dissociation limit. The I^{35}Cl(B,v′=2) products are formed with significantly less rotational excitation via this mechanism. Dynamical effects believed to be associated with tunneling through the centrifugal barrier are observed and result in highly non-statistical rotational distributions.